i am a three digit number. among my factors are 5 and 8. by 3 digits add up to only 2. what number am i?
Answer by ankor@dixie-net.com(22740)

i am a three digit number.
100a+10b+c
:
among my factors are 5 and 8.
The last digit has to be 0, a number ending in 5 cannot have a factor of 8
How about 200?
Or -240?
:
my 3 digits add up to only 2.
a + b + c = 2
-2 + 4 + 0 = 2